find the straight line represented by 2x^2-5xy+3y^2=0
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

That is not the equation of a straight line.
Cheers,
Stan H.

Answer by ikleyn(52788)

.
find the straight line represented by 2x^2-5xy+3y^2=0
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Factor the left side:

(2x-3y)*(x-y) = 0.


The equations of straight lines are

2x - 3y = 0,

and

x - y = 0.